OpenStudy (anonymous):

For a closed cylinder with radius r cm and height h cm, find the dimensions giving the minimum surface area, given that the volume is 40 cm3. Give exact answers... So V=pi * r^2 * h and A=pi * r * (2h+r) and so i set them like this? 40 = pi * r^2 * h am i doing this correctly so far?

OpenStudy (phi):

it looks like your area has only the bottom = \( \pi r^2 \) you also want the area of the top.

OpenStudy (anonymous):

ohh so A= 2πr^2 * (2h+r) ?

OpenStudy (phi):

you will want to solve for h or r using 40 = pi * r^2 * h and use that in A= 2 pi r (h+r) then set the derivative of A = 0

OpenStudy (anonymous):

okay so h= 40/ pi r^2 ? so A = 2 pi r (40/pi r^2 + r ) ?

OpenStudy (anonymous):

oh yay it's working again!!

OpenStudy (anonymous):

so is that right @phi ? h = 40/ πr^2 ?

OpenStudy (agent0smith):

This is the exact same thing as the last question, except the volume is 40 cm^3 instead of 8 cm^3

OpenStudy (anonymous):

yeah :/ so confusing!! but i figured it out :) yay!!